# Airdrops: Giving Money Away Is Harder Than It Seems

### v1

@@ 6. Design Guidelines
Target well-known, reputable entities. Instead of rewarding pseudonymous users, one can target DAOs that are building relevant apps. In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs (Foundation, 2023a). Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students (Foundation, 2023c) or members of the technical community (Foundation, 2023b) who wish to take part in research and build tools relevant to the platform. Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world (Foundation, 2021).
@@
Proactive oversight and community engagement. While the airdrop is ongoing, the platform’s data should be carefully analyzed to verify that no one is attempting to game the system. A good example was put forth by Linea, whose team uncovered a vulnerability (Liena, 2023) that allowed users to cheat their incentive mechanism. The prompt discovery prevented cheaters from obtaining over one-third [^2] A dashboard is available on Dune at https://dune.com/rusty378/guildxyz-pins. of non fungible tokens (NFTs)offered as incentives.
@@
A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities (Team, 2023), whether these are actively used or not. For example, a disclosure by an AzukiDAO community member allowed the platform to mitigate a vulnerability relatively quickly (MetaSleuth, 2023). A platform’s oversight should extend beyond on-chain data. In particular, social media is rife with scammers who advertise “fake” airdrops for their platforms, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds. Even platforms that are not planning an airdrop are targeted by such ploys.
@@
Actively engaging with the community regarding technical matters can prove beneficial, as well. For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar (cygaar, 2023), who found that it could be implemented more cost-effectively.
Starkware’s governance team have so far performed admirably in this regard (Chevet, 2023). The team organized “town hall meeting”-esque discussions and workshops, and actively invited community members to come, participate and voice their opinions. These events provide an effective and transparent communication channel, directly reaching active community members. Operating transparently and giving the community visibility into the inner-workings of the platform can assure that even if technical issues arise, they are met with understanding by users.
@@
Rewards should scale with costs. The ramifications of Goodhart’s law are readily observable in previous airdrops. For example, some airdrops attempt to incentive user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the platform in the past. Consequently, users can game the system by performing meaningless wash-trading transactions, meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the platform.
@@
This is further compounded by the fact that action-based metrics used by airdrops do not usually account for the actual costs incurred by users for each action. For instance, if the number of transactions is used as a metric, one should consider that if transaction fees are low, an airdrop farmer may cheaply perform the required amount of transactions. On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop. To remedy these issues, rewards should scale with the costs that are incurred by users.
7. Related Work
@@
The majority of recent papers on airdrops concern themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
@@
Studies. Makridis et. al. (Makridis et al., 2023) investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges. They find that airdrops of governance tokens, in particular, are positively associated with growth in market capitalization and volume.
@@
Lommers et. al. (Lommers et al., 2023) offer a broad treatment of the different kinds of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based). They offer examples of projects like Blur and the pitfalls of their airdrops, together with some advice on airdrop implementation strategies. In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success and we provide practical guidance on how to toggle these.
@@
Fan et. al.(Fan et al., 2023) study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristic of community members and the effectiveness of airdrop. They find that users receive more rewards tend to take positive actions towards the community. They summarize several arbitrage patterns and confirm the current detection is not sufficient in screening out airdrop hunters. They use an unsupervised hierarchical clustering method to capture the behaviors of eligible airdrop addresses after receiving rewards and to divide them into interpretable categories.
@@
Allen (Allen, 2023) examines airdrop design as co-evolutionary process and makes three contributions. First, applying a co-evolutionary lens to airdrops, where participants learn and adapt their strategies (e.g., new criteria, on-chain farming) both within and between web3 ecosystems. Second, nine airdrop cases studies (e.g., Optimism, Arbitrum, Blur) that demonstrate the expanding set of airdrop strategies (e.g., task-based claiming). Third, implications for airdrop design including more feedback loops to facilitate learning (e.g., multi-round airdrops), ongoing innovation in airdrop criteria (e.g., randomness), and a reversion back to simple designs for some projects (because of the costs of airdrop design).
@@
Allen et. al.(Allen et al., 2023) consider why projects might choose this unusual and costly means of token distribution. It considers a selection of high-profile airdrops as case studies between 2014 and 2022. This is the first comprehensive analysis of the rationales and mechanisms of Web3 token airdrops. We find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralisation of ownership and control of a project (building community, providing regulatory protection, and enhancing security). While many commentators describe airdrops as primarily for marketing, they claim that marketing is a weak rationale for airdropping tokens and the evidence of successful marketing-driven airdrops lacking.
@@
Technical Aspects of Airdrops. Frowis et. al. (Fröwis and Böhme, 2019) point out that efficient transfers to a large number of recipients in the case of airdrops present issues on Ethereum. They find that cost savings of a factor of two are in fact possible, but require specific provisions in the smart contract implementing the token system. Pull-based approaches, which use on-chain interaction with the recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient. Total costs are broadly linear in the number of recipients independent of the technical approach.
@@
Wahby et. al. (Wahby et al., 2020) point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities. In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely-used user credentials (ECDSA and RSA). Private airdrop for RSA builds upon a new zero-knowledge argument of knowledge of the factorization of a committed secret integer, which may be of independent interest. Their fastest implementation takes 40–180 ms to generate and 3.7–10 ms to verify an RSA private airdrop signature.
8. Conclusions
In this work we conducted analysis for of the common pitfalls of airdrops and propose concrete ways to improve the situation. Specifically, we analyze the data pertaining to four recent airdrops and conclude that the majority of proceeds (as high as 95%, typically only within a couple of transfer steps) have been sold via exchanges, implying that the airdrops fall far short of their intended purpose. We hope that the suggestions in this paper provide inspiration for researchers and practitioners alike.
Appendix A Reproducibility
To ensure complete reproducibility of our empirical analysis, we have openly shared both our dataset and scripts (Messias et al., 2023b). The primary data source employed in Section 3 comprises blockchain data extracted from both Ethereum and zkSync Era archive nodes. We meticulously collected all blocks, transactions, and events associated with the contracts of the four airdrops under study.
For Section 4, we sourced data from Growthepie.xyz (Growthepie, 2023). This platform provided various metrics (daily transaction count, daily active addresses, median transaction cost, total value locked, fees paid by users, stablecoin market cap) for several key platforms (zkSync Era, OP Mainnet, Arbitrum, Immutable X, Polygon zkEVM). All these metrics were smoothed using a 7 day moving average.
3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 6.2 Target Well-known Reputable Entities
Instead of rewarding pseudonymous users, one can target DAOs that are building relevant apps. In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs [24]. Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students [26] or members of the technical community [25] who wish to take part in research and build tools relevant to the protocol. Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world [28].
6.3 Proactive Oversight and Community Engagement
@@
While the airdrop is ongoing, the protocol’s data should be carefully analyzed to verify that no one is attempting to game the system. A good example was put forth by Linea, whose team uncovered a vulnerability [49] that allowed users to cheat their incentive mechanism. The prompt discovery prevented cheaters from obtaining over one-third [^2] A dashboard is available on Dune at https://dune.com/rusty378/guildxyz-pins. of NFTsoffered as incentives.
@@
A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities [71], whether these are actively used or not. For example, a disclosure by an AzukiDAO community member allowed the protocol to mitigate a vulnerability relatively quickly [58]. A protocol’s oversight should extend beyond on-chain data. In particular, social media is rife with scammers who advertise “fake” airdrops for their protocols, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds. Even protocols that are not planning an airdrop are targeted by such ploys.
@@
Actively engaging with the community regarding technical matters can prove beneficial, as well. For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar [16], who found that it could be implemented more cost-effectively. Operating transparently and giving the community visibility into the inner-workings of the protocol can assure that even if technical issues arise, they are met with understanding by users.
6.4 Rewards Should Scale with Costs
@@
The ramifications of Goodhart’s law are readily observable in previous airdrops. For example, some airdrops attempt to incentivize user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the protocol in the past. Consequently, users can game the system by performing meaningless wash-trading transactions [74], meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the protocol.
@@
This is further compounded by the fact that action-based metrics used by airdrops do not usually account for the actual costs incurred by users for each action. For instance, if the number of transactions is used as a metric, one should consider that if transaction fees are low, an airdrop farmer may cheaply perform the required amount of transactions. To address this, adopting a quadratic voting system could deter users from artificially inflating transaction volumes, as their rewards would be proportionate to the square root of their activities [11, 48]. Consequently, the cost to achieve a similar impact would rise, acting as a barrier against manipulation. On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop [60]. To remedy these issues, rewards should scale with the costs that are incurred by users.
6.5 Data Availability
Blockchain data is typically publicly accessible but not easily accessible. Ensuring data availability is crucial for monitoring airdrop activities and combating Sybil attacks. While previous research on airdrops and Sybil detection has primarily focused on single blockchains, our study expands this scope to multiple blockchains.
Specifically, we collect data from two major Ethereum rollups, Arbitrum and zkSync Era, and gather and label data from significant airdrops, including a major Sybil attack on zkSync Era (the Gemstone “airdrop”). Since Sybil detection is increasingly challenging and requires data from multiple chains—given that protocols often reward participants from different chains [27, 4]—data availability is essential for analyzing past airdrops and testing new metrics for future applications. However, the rapid growth of blockchain data can make its collection and processing impractical.
Protocols should ensure their data is available for researchers and companies, helping improve monitoring tools and facilitate learning from the data. To support this, we will make our zkSync Era and Ethereum datasets publicly available, enabling others to utilize them for optimizing Sybil detection and airdrop designs.
7 Related Work
@@
The majority of recent papers on airdrops concerned themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
@@
Studies.  Makridis et. al. [53] investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges. They find that airdrops of governance tokens, in particular, are positively associated with growth in market capitalization and volume.
@@
Lommers et. al. [52] offer a broad treatment of the different types of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based). They offer examples of projects like Blur and the pitfalls of their airdrops, along with some advice on airdrop implementation strategies. In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success, and they provide practical guidance on how to toggle these.
@@
Fan et. al.[21] study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristics of community members and the effectiveness of airdrop. They find that users who receive more rewards tend to take positive actions towards the community. They summarize several arbitrage patterns and confirm that the current detection is not sufficient to detect airdrop hunters. They use an unsupervised hierarchical clustering method to capture the behaviors of eligible airdrop addresses after receiving rewards and to divide them into interpretable categories. Other works focused on addressing the problem of Sybil detection, providing valuable insights into identifying airdrop farmers. Commonly adopted practices include graph network analysis [21, 51] and machine learning techniques [36, 79, 73, 13, 59]. The former involves identifying prevalent transfer patterns, such as sequences, radial structures, or a combination of both. The latter utilizes approaches like node embedding and employs universal pre-trained transformer encoders like BERT for effective fraud detection [36] or clustering [73, 59].
@@
Allen [2] examines airdrop design as co-evolutionary process and makes three contributions. First, applying a co-evolutionary lens to airdrops, where participants learn and adapt their strategies (e.g., new criteria, on-chain farming) both within and between web3 ecosystems. Second, nine airdrop cases studies (e.g., Optimism, Arbitrum, Blur) that demonstrate the expanding set of airdrop strategies (e.g., task-based claiming). Third, implications for airdrop design including more feedback loops to facilitate learning (e.g., multi-round airdrops), ongoing innovation in airdrop criteria (e.g., randomness), and a reversion back to simple designs for some projects (because of the costs of airdrop design).
@@
Allen et. al.[3] consider why projects might choose this unusual and costly means of token distribution. It considers a selection of high-profile airdrops as case studies between 2014 and 2022. This is the first comprehensive analysis of the rationales and mechanisms of Web3 token airdrops. They find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralization of ownership and control of a project (building community, providing regulatory protection, and enhancing security). While many commentators describe airdrops as primarily for marketing, they claim that marketing is a weak rationale for airdropping tokens and the evidence of successful marketing-driven airdrops lacking.
@@
Technical Aspects of Airdrops  Frowis et. al. [31] point out that efficient transfers to a large number of recipients in the case of airdrops present problems on Ethereum. They find that cost savings of a factor of two are in fact possible, but require specific provisions in the smart contract implementing the token system. Pull-based approaches, which use on-chain interaction with recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient. Total costs are broadly linear in the number of recipients, regardless of the technical approach.
@@
Wahby et. al. [75] point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities. In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely used user credentials (ECDSA and RSA). Private airdrop for RSA builds upon a new zero-knowledge argument of knowledge of the factorization of a committed secret integer, which may be of independent interest. Their fastest implementation takes 40–180 ms to generate and 3.7–10 ms to verify an RSA private airdrop signature.
8 Conclusions
In this study, we investigated the prevalent issues with airdrops and developed guidelines to improve their impact. Our study analyzed six notable airdrops, uncovering a pattern in which recipients quickly traded their airdropped tokens on exchanges. Specifically, 36.51% of ENS tokens, 35.35% of dYdX tokens, and 53.47% of 1inch tokens were exchanged soon after distribution. Generally, these tokens were sold after only 1 or 2 transactions, with a median of 2 transactions, demonstrating that airdrops do not promote sustained engagement or attract contributors of significant value. Furthermore, our examination of Arbitrum indicated a significant increase in total daily fees during the airdrop event, but a decrease in transaction count per address afterward. In particular, other protocols outperformed Arbitrum in the absence of an airdrop. The average transaction fees across all protocols followed a similar trend since June 2023, suggesting that airdrops may not be the primary driver of user base growth and can impact activity on other chains.
We also acknowledge the common airdrop challenges related to airdrop farmers, governance token distribution, and insider trading. These challenges serve as valuable lessons for blockchain protocols considering airdrop tokens and offer insight into the potential path to success. As a measure to avoid these pitfalls, we propose design guidelines for airdrop campaigns, which are summarized below.
Alternative incentives may better sustain user engagement.  An interesting approach involves offering alternative incentives, such as fee discounts for future interactions within the protocol, to foster long-term engagement. Additionally, implementing multi-round airdrops allows for learning from past rounds, enabling iterative improvements to address previously identified issues. Another approach involves exploring innovative metrics to assess the eligibility of an address for receiving an airdrop. This might also be a promising tactic to address concerns related to airdrop farming who typically tries to game the system by utilizing past airdrop actions. However, this approach may not completely solve the challenge of insider trading.
Target well-known reputable entities.  Prioritizing established and reputable entities, including those building on the protocol, research groups, tech communities, and students, can contribute to sustained engagement. Funding these entities or directing airdrops towards DAOs developing relevant dApps presents an opportunity to attract value-driven users and promote long-term involvement.
Proactive oversight and community engagement.  Critical to the success of any airdrop is proactive oversight and community engagement. Analyzing data from past airdrops, even those on different chains, provides insights into pitfalls and allows for adjustments in future airdrop strategies. Open communication channels with the community help understand their interests and desired protocol functionalities.
Rewards should scale with costs.  Ensure that rewards scale proportionally with associated costs to prevent exploitation by airdrop farmers. Aligning rewards with costs is a strategic measure to mitigate potential gaming of the system.
In conclusion, our paper offers insights into airdrop challenges and proposes practical guidelines for designing effective airdrop campaigns. We hope these suggestions inspire blockchain protocols, researchers, and practitioners. Finally, in our commitment to promoting scientific reproducibility, we plan to share the dataset and scripts utilized in this study [57].